OpenStudy (anonymous):

If ST = 21, SP = 3b – 11, and PT = b + 4, solve for b.

OpenStudy (gorv):

SP-ST=PT PT=21-3B-11=10-3B

OpenStudy (gorv):

but pt =b+4 b+4=10-3b solve it further noe itzz easy